Can someone help me with this question, please?
One of the roots of the equation 4x^2-2px+p=0, p is not = 0 is two times the other root. Find the possible value of p.

Answer by venugopalramana(3286)

One of the roots of the equation 4x^2-2px+p=0, p is not = 0 is two times the other root. Find the possible value of p.
LET ONE ROOT BE =K..THEN THE OTHER ROOT IS 2K
HENCE SUM OF ROOTS =K+2K=3K= -(-2P)/4=P/2....OR.... K=P/6..................I
PRODUCT OF THE ROOTS =K*2K=2K^2=P/4............................II
SUBSTITUTING I IN II...WE GET
2*(P/6)^2=P/4
2P^2/36=P/4......CROSS MULTIPLYING
8P^2=36P
2P^2=9P
2P^2-9P=0
P(2P-9)=0
SINCE P IS NOT EQUAL TO 0 ,WE GET
2P-9=0
2P=9
P=9/2